Evaluate the following line integrals along with the curve C.

(a) ∫

C (x2 − 2y2 ) ds; C is the line segment parametrized by r(t) = < t / √ 2 , t / √ 2 >, for 0 ≤ t ≤ 4.


1
Expert's answer
2021-12-29T13:00:19-0500

"\int_C(x^2 − 2y^2 ) ds"

"r(t) = < \dfrac{t}{\sqrt{2}} , \dfrac{t}{\sqrt{2}} >, 0 ≤ t ≤ 4."


"x^2 − 2y^2=\dfrac{t^2}{2}-t^2=-\dfrac{t^2}{2}"

"\dfrac{dx}{dt}=\dfrac{1}{\sqrt{2}}, \dfrac{dy}{dt}=\dfrac{1}{\sqrt{2}}"

"\sqrt{(\dfrac{dx}{dt})^2+(\dfrac{dy}{dt})^2}=\sqrt{(\dfrac{1}{\sqrt{2}})^2+(\dfrac{1}{\sqrt{2}})^2}=1"

"\int_C(x^2 − 2y^2 ) ds=\displaystyle\int_{0}^{4}(-\dfrac{t^2}{2})(1)dt"

"=\big[-\dfrac{t^3}{6}\big]\begin{matrix}
 4 \\
 0
\end{matrix}=-\dfrac{32}{3}"